# Distributional Adversarial Attacks and Training in Deep Hedging

### v1

@@ Abstract
In this paper, we study the robustness of classical deep hedging strategies under distributional shifts by leveraging the concept of adversarial attacks. We first demonstrate that standard deep hedging models are highly vulnerable to small perturbations in the input distribution, resulting in significant performance degradation. Motivated by this, we propose an adversarial training framework tailored to increase the robustness of deep hedging strategies. Our approach extends pointwise adversarial attacks to the distributional setting and introduces a computationally tractable reformulation of the adversarial optimization problem over a Wasserstein ball. This enables the efficient training of hedging strategies that are resilient to distributional perturbations.
Through extensive numerical experiments, we show that adversarially trained deep hedging strategies consistently outperform their classical counterparts in terms of out-of-sample performance and resilience to model misspecification. Our findings establish a practical and effective framework for robust deep hedging under realistic market uncertainties.
The code is available on https://github.com/Guangyi-Mira/Distributional-Adversarial-Attacks-and-Training-in-Deep-Hedging.
@@ 1 Introduction
The concept of robustness is important across machine learning applications. In image classification tasks, szegedy2014intriguing first highlighted that neural networks are vulnerable to adversarial attacks—small, imperceptible changes to input images that cause misclassification. This discovery led to a growing body of research on adversarial attacks and training techniques to improve network robustness against them (goodfellow2014explaining ; madry2018towards ; zhang2019theoretically ). However, most studies have focused on pointwise adversarial perturbations, where attacks modify individual data points. A broader perspective considers distributional attacks, where an adversary perturbs the entire data-generating distribution rather than specific samples (bai2023wasserstein ; dong2024enhancing ; ref:Kuhn:tutorial-19 ; ref:Kuhn:DRObook-24 ; staib2017distributionally ; ref:Gao-23 ). Specifically, the distributional version of adversarial attacks can be applied to search for the worst-case scenario in the DRO problem.
@@
  • We propose the first framework that unifies distributional adversarial training with deep hedging. Our approach combines tractable reformulation techniques from Wasserstein DRO staib2017distributionally , with sensitivity analysis results bai2023wasserstein , to derive a computationally efficient approximation of adversarial attacks within a Wasserstein ball. This provides both a theoretical foundation and practical algorithms for robust decision-making.
  • We demonstrate that classical deep hedging strategies are vulnerable to distributional adversarial attacks, where even minor input distribution shifts can cause substantial performance degradation. To address this, we propose an adversarial training framework that iteratively generates perturbed adversarial distributions from training data and optimizes the hedging model against these scenarios. This method provides a computationally tractable solution for Wasserstein DRO in deep hedging.
@@
Further related work.
In addition to the works wu2023robust ; lutkebohmert2022robust ; limmer2024robust discussed earlier, since its introduction in buehler2019deep numerous works have studied deep hedging and related machine learning-based hedging methods from different perspectives. Examples include tackling complex pricing problems carbonneau2021equal ; carbonneau2021deep , targeting improved efficiency imaki2021no ; Mueller2024 or general initial portfolios murray_2022_deep , alternative reinforcement learning frameworks kolm2019dynamic ; du2020deep ; cao2021deep ; vittori2020option , empirical approaches ruf2022hedging ; mikkila2023empirical , or implementations on quantum hardware raj2023quantum . We refer to the surveys hambly_2023_recen ; ruf_2020_neura for a comprehensive overview.
@@ 2 Deep Hedging Framework
We consider a Black-Scholes model with entropic risk and a Heston model with CVaR. Detailed formulations of both models and the associated loss functions are provided in the appendix.
@@ 3 Adversarial Attacks
Equation (C.5) provides an approximate worst-case perturbation by leveraging the sensitivity expansion of the Wasserstein DRO problem, which directly yields a computationally tractable solution that approximates (3.4). However, this approach may not fully explore the adversarial landscape. Therefore, we instead constrain optimization to a subset of the ambiguity set which contains $\eta_{\delta}$ and has a simpler structure. The set $\hat{B}_{\delta}(\mu_{N})$ is defined as
@@ 4 Algorithms for Distributional Adversarial Attacks in Deep Hedging
We first focus on the Black-Scholes model, where the input information is exactly the single price trajectory S, as detailed in the appendix. We will then explain how to extend the algorithm to models with more than one trajectory, including the Heston model, where the input is a pair of price and variance trajectories S and v. For the readers’ convenience, we focus on these two cases here and present algorithms for a general setting in the Appendix.
@@
The infinity norm has the $L^{1}$-norm as its dual norm, i.e., $\|g\|_{*}=\sum\nolimits_{t=0}^{T}|g_{t}|$ and $h(x)=\text{sign}(x)$, applied component-wise, satisfy $\langle h(x),x\rangle=\|x\|_{*}$, which implies Assumption 3.1.
@@
Following the above setup, equation (C.5) provides an asymptotic approximation of the worst-case distributional shift, leading to a one-step perturbation of each path like FGSM
@@
where $\text{dist}=(\frac{1}{N}\sum\nolimits_{i=1}^{N}d(\textbf{S}_{i},\hat{\textbf{S}}_{i})^{p})^{1/p}$.
The overall procedure, called the Wasserstein Projection Gradient Descent (WPGD), is summarized in Algorithm 1.
@@
10:  $\hat{\textbf{S}}_{1},\dots,\hat{\textbf{S}}_{n}\leftarrow\text{Proj}_{\delta}(\hat{\textbf{S}}_{1},\dots,\hat{\textbf{S}}_{n})$
@@ 5.1 Attacks on classical deep hedging strategies
To understand practical implications and quantify the extent of distortion in the path space due to the attack, we assessed perturbation impacts through covariance matrix comparisons, as shown in Table 2. The Frobenius distance between the covariance matrices of perturbed and original paths, given the original covariance matrix norm ($\approx 386$), indicates minor covariance distortions, especially for small perturbations ($\delta<0.1$). Specifically, simultaneous perturbations on both price and variance processes (SV-Attack) result in less pronounced covariance changes compared to perturbations only on the price process (S-Attack). In the appendix we also report autocorrelation function comparisons.
Overall, together with with Table 1 these analyses show that adversarial attacks may significantly deteriorate the hedging strategy’s performance, even under seemingly modest perturbations, thereby underscoring the necessity for robust neural network models in financial applications.
@@ 5.2 Adversarial training
We now introduce adversarial training, which aims to enhance the robustness of strategies against distributional adversarial attacks and hence solve the DRO problem. We adopt the standard deep hedging methodology buehler2019deep as baseline and expand it to incorporate adversarial examples during training. Full experimental details (e.g., network architecture, hyperparameters) are provided in the appendix.
@@
We conduct our experiments using two well-established financial models: the Black-Scholes and the Heston model (detailed in the appendix). Here we report results for the Heston model. Results for the BS model are very similar (see appendix). For each model, we generate extensive training datasets of 100,000 sample paths. To examine robustness across varying dataset sizes, we partition each dataset into smaller subsets with sizes $N$ ranging from 5,000 to 100,000 samples. Neural networks are independently trained on these subsets, and the average performance is assessed and reported on a fixed test set, which contains 1 million paths, generated from the same distribution. In addition, we generate a validation set of 100,000 paths, but only $N$ paths will be used for validation, so that the training is exposed to only a limited number of data depending on $N$.
5.3 Robustness of adversarial training
For the robust strategy trained using adversarial training detailed in Section 5.2, in this section we first examine the adversarial loss to confirm training effectiveness. We evaluate robust strategies trained under distributional adversarial attacks with perturbation magnitudes $\delta=0.01$ and $0.1$, comparing their performance against a clean strategy trained on an identical dataset of 100,000 samples. Table 3 illustrates the adversarial loss for each strategy under varying perturbation levels. The robust strategies consistently yield significantly lower adversarial losses compared to the clean strategy, clearly demonstrating that our robust training framework successfully provides a computationally tractable approximation of the theoretically optimal distributionally robust optimization (DRO) solution, enhancing model resilience against adversarial perturbations. Additionally, there is an expected trade-off observed in the loss: for the robust strategy trained with $\delta=0.1$, the loss at no or small perturbation levels is higher, but the adversarial loss at larger perturbation magnitudes is substantially reduced, aligning with the designed training objective.
$\delta$ 0 0.01 0.03 0.05 0.1 0.3 0.5
clean 1.9162 1.9598 2.0530 2.1560 2.4659 4.6069 8.0379
@@ 5.4 Adversarial training improves out-of-sample and out-of-distribution performance
6 Numerical Experiment on Real Market Data
In this section, we evaluate the performance of adversarial training on real market data. Specifically, we train hedging strategies using historical daily closing prices from leading companies in the S&P 500 index from yfinance , covering the period from 26 September 2008 to 8 March 2020. The time period is chosen in line with the benchmark method lutkebohmert2022robust with which we compare our method.
For this evaluation, we constructed two synthetic datasets based on an additional model introduced in the appendix.
  • The FIX dataset is simulated using the General Affine Diffusion (GAD) model with fixed parameters estimated from a 250-day period prior to 8 March 2020.
  • The ROBUST dataset, as proposed by lutkebohmert2022robust , is also based on the GAD model. It incorporates parameter robustness by sampling parameters uniformly from intervals. These intervals are constructed using the extreme values across 26 parameter estimates, which are obtained from rolling 250-day windows updated every 100 days between 26 September 2008 and 8 March 2020.
For each synthetic dataset (FIX and ROBUST), we generate 100,000 paths each for training, validating, and testing. The starting price is set as the closing price of the respective company on 8 March 2020, and each trajectory is then scaled to begin from 10.
Directly training on the FIX and ROBUST datasets corresponds to the methods proposed in Deep Hedging buehler2019deep and Robust Deep Hedging lutkebohmert2022robust , respectively. Building on this foundation, we implement the adversarial framework described in Section 5.2, resulting in two adversarial variants of the strategies. Similar to lutkebohmert2022robust , we evaluate out-of-sample performance on real data by computing profit and loss (PnL) for hedging strategies on the tested price trajectories starting form 9 March 2020. Table 4 summarizes results for five companies: Apple (AAPL), Amazon (AMZN), Microsoft (MSFT), Google (GOOGL), and Berkshire Hathaway (BRK-B). To address the limitations of using a single price trajectory, including the absence of validation data and the training randomness, we adopt the following evaluation protocol. For adversarial training, we use N=100, [^0] N=100{,}000 training samples and report the average out-of-sample performance across six strategies corresponding to $(\alpha,\delta)\in\{0,1\}\times\{0.03,0.05,0.1\}$. For clean training, we average the results over three independent runs with identical settings. Bold entries indicate top-performing results per company.

### v2

@@ Abstract
In this paper, we study the robustness of classical deep hedging strategies under distributional shifts by leveraging the concept of adversarial attacks. We first demonstrate that standard deep hedging models are highly vulnerable to small perturbations in the input distribution, resulting in significant performance degradation. Motivated by this, we propose an adversarial training framework tailored to increase the robustness of deep hedging strategies. Our approach extends pointwise adversarial attacks to the distributional setting and introduces a computationally tractable reformulation of the adversarial optimization problem over a Wasserstein ball. This enables the efficient training of hedging strategies that are resilient to distributional perturbations. Through extensive numerical experiments, we show that adversarially trained deep hedging strategies consistently outperform their classical counterparts in terms of out-of-sample performance and resilience to model misspecification. Additional results indicate that the robust strategies maintain reliable performance on real market data and remain effective during periods of market change. Our findings establish a practical and effective framework for robust deep hedging under realistic market uncertainties. [^1] The code is available on https://github.com/Guangyi-Mira/Distributional-Adversarial-Attacks-and-Training-in-Deep-Hedging
@@ 1 Introduction
The concept of robustness is important across machine learning applications. In image classification tasks, szegedy2014intriguing first highlighted that neural networks are vulnerable to adversarial attacks—small, imperceptible changes to input images that cause misclassification. This discovery led to a growing body of research on adversarial attacks and training techniques to improve network robustness against them (goodfellow2014explaining ; madry2018towards ; zhang2019theoretically ). However, most studies have focused on pointwise adversarial perturbations, where attacks modify individual data points. A broader perspective considers distributional attacks, where an adversary perturbs the entire data-generating distribution rather than specific samples (bai2023wasserstein ; bai2025wasserstein ; dong2024enhancing ; ref:Kuhn:tutorial-19 ; ref:Kuhn:DRObook-24 ; staib2017distributionally ; ref:Gao-23 ). Specifically, the distributional version of adversarial attacks can be applied to search for the worst-case scenario in the DRO problem.
@@
  • We propose the first framework that unifies distributional adversarial training with deep hedging. Building on tractable reformulations from Wasserstein DRO staib2017distributionally and sensitivity analysis results bai2023wasserstein , our approach derives computationally efficient adversarial attacks within a Wasserstein ball and integrates them into an adversarial training procedure. This framework reveals the vulnerability of classical deep hedging to even small distributional shifts and provides a tractable method to enhance robustness.
@@
  • We validate our framework on real market data, demonstrating that adversarially trained strategies remain effective and robust during periods of market stress and distributional shifts, and compare our results against the robust deep hedging approach of lutkebohmert2022robust .
Related work.
Deep hedging was introduced in buehler2019deep ; buehler2019deep2 . Since its introduction, numerous works have studied deep hedging and related machine learning-based hedging methods from different perspectives. Robustness of deep hedging methods has been studied in wu2023robust ; lutkebohmert2022robust ; limmer2024robust . Further directions include tackling complex pricing problems carbonneau2021equal ; carbonneau2021deep ; biagini2023 , targeting improved efficiency imaki2021no ; Mueller2024 or general initial portfolios murray_2022_deep , alternative reinforcement learning frameworks kolm2019dynamic ; du2020deep ; cao2021deep ; vittori2020option , empirical approaches ruf2022hedging ; mikkila2023empirical , or implementations on quantum hardware raj2023quantum . We refer to the surveys hambly_2023_recen ; ruf_2020_neura for a comprehensive overview.
Parallel to this, distributionally robust optimization (DRO) has emerged as a framework for decision-making under model uncertainty, with Wasserstein-based methods offering both theoretical guarantees and computational tractability ref:Kuhn:DRObook-24 ; ref:Peyman-18 ; ref:Kuhn:tutorial-19 . In machine learning, adversarial training provides another robustness perspective, where worst-case perturbations improve generalization szegedy2014intriguing ; goodfellow2014explaining ; madry2018towards ; zhang2019theoretically . Recent works extend pointwise adversarial attacks to the distributional setting bai2023wasserstein ; bai2025wasserstein ; dong2024enhancing ; staib2017distributionally ; ref:Gao-23 , thereby connecting adversarial training methods with DRO formulations.
@@ 2 Deep Hedging Framework
We consider a Black-Scholes model and General Affine Diffusion model with entropic risk and a Heston model with CVaR. Detailed formulations of these models and the associated loss functions are provided in the appendix.
@@ 3 Adversarial Attacks
Equation (3.9) provides an approximate worst-case perturbation by leveraging the sensitivity expansion of the Wasserstein DRO problem, which directly yields a computationally tractable solution that approximates (3.4). However, this approach may not fully explore the adversarial landscape. Therefore, we instead constrain optimization to a subset of the ambiguity set which contains $\eta_{\delta}$ and has a simpler structure. The set $\hat{B}_{\delta}(\mu_{N})$ is defined as
@@ 4 Algorithms for Distributional Adversarial Attacks in Deep Hedging
We first focus on the Black-Scholes model, where the input information is exactly the single price trajectory S, as detailed in the appendix. We will then explain how to extend the algorithm to models with more than one trajectory, including the Heston model, where the input is a pair of price and variance trajectories S and v. For the readers’ convenience, we focus on these two cases here and present algorithms for a general setting in the supplementary material.
@@
The infinity norm has the $L^{1}$-norm as its dual norm, i.e., $\|g\|_{*}=\sum\nolimits_{t=0}^{T}|g_{t}|$, and $h(x)=\text{sign}(x)$, applied component-wise, satisfies $\langle h(x),x\rangle=\|x\|_{*}$, which implies Assumption 3.1.
@@
Following the above setup, equation (3.9) provides an asymptotic approximation of the worst-case distributional shift, leading to a one-step perturbation of each path like FGSM
@@
where $\text{dist}=(\frac{1}{N}\sum\nolimits_{i=1}^{N}d(\textbf{S}_{i},\hat{\textbf{S}}_{i})^{p})^{1/p}$. The overall procedure, called the Wasserstein Projection Gradient Descent (WPGD), is summarized in Algorithm 1.
@@
10:  $\hat{\textbf{S}}_{1},\dots,\hat{\textbf{S}}_{n}\leftarrow\text{Proj}_{\hat{B}_{\delta}(\mu)}(\hat{\textbf{S}}_{1},\dots,\hat{\textbf{S}}_{n})$
@@ 5.1 Attacks on classical deep hedging strategies
To understand practical implications and quantify the extent of distortion in the path space due to the attack, we assessed perturbation impacts through covariance matrix comparisons, as shown in Table 2. The Frobenius distance between the covariance matrices of perturbed and original paths, given the original covariance matrix norm ($\approx 386$), indicates minor covariance distortions, especially for small perturbations ($\delta<0.1$). Specifically, simultaneous perturbations on both price and variance processes (SV-Attack) result in less pronounced covariance changes compared to perturbations only on the price process (S-Attack). In the appendix we also report autocorrelation function comparisons. Overall, together with with Table 1, these analyses show that adversarial attacks may significantly deteriorate the hedging strategy’s performance, even under seemingly modest perturbations, thereby underscoring the necessity for robust neural network models in financial applications.
@@ 5.2 Adversarial training
We introduce adversarial training, which aims to enhance the robustness of strategies against distributional adversarial attacks and hence solve the DRO problem. We adopt the standard deep hedging methodology buehler2019deep as baseline and expand it to incorporate adversarial examples during training. Experimental details (e.g., network architecture, hyperparameters) are provided in the appendix.
@@
We conduct our experiments using three well-established financial models: Black–Scholes, Heston, and the General Affine Diffusion (GAD) model. Here we present results for the Heston model in Section 5.3, while model details and results for the other models are provided in the appendix. Results for real market data are reported in Section 5.4 below. For each model, we generate extensive training datasets of 100,000 sample paths. To examine robustness across varying dataset sizes, we partition each dataset into smaller subsets with sizes $N$ ranging from 5,000 to 100,000 samples. Neural networks are independently trained on these subsets, and the average performance is assessed and reported on a fixed test set, which contains 1 million paths, generated from the same distribution. In addition, we generate a validation set of 100,000 paths, but only $N$ paths will be used for validation, so that the training is exposed to only a limited number of data depending on $N$.
Robustness of adversarial training.
Strategies trained with our adversarial procedure achieve lower losses under distributional perturbations than classical deep hedging strategies, indicating that the method effectively approximates the desired distributionally robust solution. Detailed experiments and results are provided in the appendix.
@@ 5.3 Adversarial training improves out-of-sample and out-of-distribution performance
5.4 Experiments on Market Data
In this section, we evaluate the performance of adversarial training on real market data. Specifically, we train hedging strategies using historical daily closing prices from leading companies in the S&P 500 index from yfinance , covering the period from 26 September 2008 to 8 March 2020.
The time period is chosen in line with the benchmark method lutkebohmert2022robust with which we compare our method.
For this evaluation, we constructed two synthetic datasets based on an additional model introduced in the appendix. The FIX dataset is simulated using the General Affine Diffusion (GAD) model with fixed parameters estimated from a 250-day period prior to 8 March 2020. The ROBUST dataset, following lutkebohmert2022robust , is also based on the GAD model but incorporates parameter robustness by sampling parameters uniformly from intervals determined by the extreme values across 26 rolling estimates. These estimates are obtained from 250-day windows updated every 100 days between 26 September 2008 and 8 March 2020. For each dataset (FIX and ROBUST), we generate 100,000 paths for training, validation, and testing. The starting price is set as the closing price of the respective company on 8 March 2020, and each trajectory is scaled to begin from 10.
Directly training on the FIX and ROBUST datasets corresponds to the methods proposed in Deep Hedging buehler2019deep and Robust Deep Hedging lutkebohmert2022robust , respectively. Building on this foundation, we implement the adversarial framework described in Section 5.2, resulting in two adversarial variants of the strategies. Similar to lutkebohmert2022robust , we evaluate out-of-sample performance on real data by computing profit and loss (PnL) for hedging strategies on the tested price trajectories starting form 9 March 2020. Table 3 summarizes results for five companies. [^2] Including Apple (AAPL), Amazon (AMZN), Microsoft (MSFT), Google (GOOGL), and Berkshire Hathaway (BRK-B), where bold entries indicate top-performing results per company. To address the limitations of using a single price trajectory, we refer the reader to the Appendix for details of the evaluation protocol and an additional evaluation result.
@@ 5.4 Experiments on Market Data
6 Conclusion
We presented a robust deep hedging framework that leverages adversarial training under Wasserstein ambiguity to address the challenges of model misspecification and distributional shifts. By formulating the distributionally robust optimization problem as a minimax objective and approximating it through a tractable reformulation, we demonstrated how deep hedging strategies can be trained adversarially. Empirical results across a variety of widely used synthetic models and data regimes show that adversarially trained strategies achieve improved out-of-sample and out-of-distribution performance, especially under structural changes or limited data availability. Further experiments on real market data suggest that these strategies also generalize well beyond simulated settings, maintaining robustness in periods of market stress. These findings underscore the practical value of robust training methods in financial environments characterized by uncertainty and instability.